A charge of 23.0 nC is placed in a uniform electric field that is directed vertically upward and that has a magnitude of 4.00x104 N/C What work is done by the electric force when the charge moves
media%2F85a%2F85a31b86-dc64-4d3d-a007-c2
Answer #1

part C)

q = 23 nC = 23 *10^-9 C

E = 4 *10^4 N/C

d = 2.4 m

theta = 45 degree

work done by the electric force = -q *E * d * cos(45 degree)

work done by the electric force = -23 * 10^-9 * 4 *10^4 * 2.4 * cos(45 degree)

work done by the electric force = -0.00156 J

the work done by the electric force is -0.00156 J